OpenStudy (anonymous):

The charge by mass ratio of an ion M+ is 4.2 * 10^6 C /kg The ion may be?

OpenStudy (anonymous):

can u tell me the charge of one proton in COLOUMB

OpenStudy (anonymous):

1.6 *10^19 C

OpenStudy (anonymous):

1.6*10^(-19)C....... I guess

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

i forgot to put that

OpenStudy (anonymous):

so, 1.6*10^(-19)C / Mass = 4.2 * 10^6 C /kg MASS=3.81*10^(-26) kg

OpenStudy (anonymous):

so..which is that ion...?

OpenStudy (anonymous):

can u convert that into Kg into moles?

OpenStudy (anonymous):

Yup .... it is 22.9 gm and it is Na + thxxxx

OpenStudy (anonymous):

Welcomx